Seminar Tutorial  E.A. Gonzalez-Solares

Seminar Tutorial

Eduardo A. Gonzalez-Solares

[LightCyan] Second Slide First item Second item with a link to MY HOMEPAGE.

(\frac{dN}{dz} ) = \frac{3N Ω_{s}}{2z_{c}^{3}} z^{2} exp [− (\frac{z}{z_{c}} )^{3/2}]